A 234 ft stick is driven 26 ft into the bottom of a lake that is 89 ft deep. On the stick is a crack 41 ft below the surface of the water. A hook hangs 5 ft from the top of the pole. How far is the hook from the crack?
1
Expert's answer
2010-07-09T12:46:15-0400
The length of the stick is 234 ft. A hook hangs 5 ft from the top of the pole. To find the distance between the hook and the crack we have to find the distance between the top of the stick and the surface of the water: 234-(26+89)=119 ft, so the distance from the hook to the surface of the water is 119-5=114 ft. And the crack on the stick is 41 ft below the surface of the water. So, to find how far the hook from the crack we have to calculate 114+41=155 ft